Directions: find (f-g) (x) using the two functions given in each number.



1. F(x) = 3x+3


g(x) = -4x+1



2. F(x) = 2x+5


g(x) = 4x² +2x-2



3.f(x)=5x+1


g(x) =3x-2



4.f(x) -15 x² -2x+5


g(x) = 3x2+x-7



5.f(x) = 3x² - 2x+1


g(x) = 4x² +5x-4


1
Expert's answer
2022-03-17T18:21:41-0400

1) "(f-g)(x)=3x+3-(-4x+1)=7x+2."

2) "(f-g)(x)=2x+5-(4x^2+2x-2)=-4x^2+7."

3) "(f-g)(x)=5x+1-(3x-2)=2x+3."

4) "(f-g)(x)=-15x^2-2x+5-(3x^2+x-7)=-18x^2-3x+12."

5) "(f-g)(x)=3x^2-2x+1-(4x^2+5x-4)=-x^2-7x+5."